# Agentic Information Retrieval

### v1

@@ 3.1 Life Assistant
Agentic IR capabilities are already present in major products like Apple’s ecosystem, where Apple Intelligence [^1] https://www.apple.com/apple-intelligence/ powers advanced assistant features across devices such as iPhone, iPad, and Mac (Apple, 2024). Apple Intelligence enhances user experience by seamlessly integrating with apps, services, and smart devices, embodying the proactive and contextual characteristics of agentic IR. Beyond Apple, other life assistants, such as Google Assistant [^2] https://assistant.google.com/, Amazon Alexa [^3] https://www.alexa.com/, Oppo Breeno, and Huawei Celia [^4] https://consumer.huawei.com/en/emui/celia/, operate across diverse platforms, including smartphones, smart home devices, and wearables. These assistants empower users with convenient control over both digital and physical environments, enabling them to make informed plans and adjustments anytime, anywhere (Li et al., 2024b).
@@ 3.2 Business Assistant
The business assistant workflow generally involves four stages: query understanding, document retrieval, information integration, and response generation, as illustrated in Figure 3. Today, there are already several business assistants powered by agentic IR in use, such as Amazon Q Business [^5] https://aws.amazon.com/q/business/. Below, we will illustrate each stage with reference to this example.
@@
Query understanding. Given a business-related query, the agent, core of the business assistant, first attempts to understand and analyze the user’s intention. For complex queries, the agent can generate thoughts (Tht) with CoT to break down the problem into smaller, manageable steps, allowing for multi-step reasoning. In Amazon Q Business, conversations are continuous, allowing the historical dialogues to serve as memory (Mem), helping the agent better understand the context and user intention.
@@
Information integration. In many cases, the retrieved information is scattered across multiple sections or even across different documents. Before responding to the query, the agent generates thoughts (Tht) or uses tools (Tool) to combine and condense the information. In Amazon Q Business, the system is equipped with the RAG framework by default and supports various plugins to generate responses or complete tasks.
Response generation. Finally, the agent generates a response, reaching the resulting information state. Depending on the query and the retrieved data, the resulting information state can take various forms. The response may be presented in multiple formats, including plain text, tables, visualized charts, etc. The assistant can also complete tasks and return an action state. In Amazon Q Business, it also links the answer back to its source documents to ensure transparency, allowing users to trace how the response was derived.
@@ Acknowledgments
We thank the insightful discussions with Jie Fu, Jun Wang, Ying Wen, Zheng Tian, Minjie Wang, David Wipf, Quan Gan, Qiuying Peng, and Grace Hui Yang.

### v2

@@ 3.1 Life Assistant
Agentic IR capabilities are already present in major products like Apple’s ecosystem, where Apple Intelligence [^1] https://www.apple.com/apple-intelligence/ powers advanced assistant features across devices such as iPhone, iPad, and Mac (Apple, 2024). Apple Intelligence enhances user experience by seamlessly integrating with apps, services, and smart devices, embodying the proactive and contextual characteristics of agentic IR. Beyond Apple, other life assistants, such as Google Assistant [^2] https://assistant.google.com/, Oppo Breeno, and Huawei Celia [^3] https://consumer.huawei.com/en/emui/celia/, operate across diverse platforms, including smartphones, smart home devices, and wearables. These assistants empower users with convenient control over both digital and physical environments, enabling them to make informed plans and adjustments anytime, anywhere (Li et al., 2024b).
@@ 3.2 Business Assistant
Currently, there are already several business assistants powered by agentic IR in use, such as Microsoft 365 Copilot [^4] https://www.microsoft.com/en-us/microsoft-365/copilot/copilot-for-work, Notion AI  [^5] https://www.notion.so/product/ai, and IBM watsonx [^6] https://www.ibm.com/watsonx. Typically, the workflow of business assistant consists of four stages: query understanding, document retrieval, information integration, and response generation, as illustrated in Figure 3. Each stage is described in detail below.
@@
Query understanding. Given a business-related query, the agent, core of the business assistant, first attempts to understand and analyze the user’s intention. For complex queries, the agent can generate thoughts (Tht) with CoT to break down the problem into smaller, manageable steps, allowing for multi-step reasoning. In business assistants, conversations are continuous, allowing the historical dialogues to serve as memory (Mem), helping the agent better understand the context and user intention.
@@
Information integration. In many cases, the retrieved information is scattered across multiple sections or even different documents. To construct a comprehensive response, the agent must combine and distill the information, which often requires functions like Tht and Tool. By generating internal thoughts (Tht), the agent can establish logical connections between disparate pieces of information and perform complex reasoning, gradually working toward the resulting information state. Additionally, specialized tools (Tool) assist in this process by enabling capabilities such as executing mathematical calculations and filtering out redundant information.
Response generation. Finally, the agent generates a response, reaching the resulting information state. Depending on the query and the retrieved data, the resulting information state can take various forms. The response may be presented in multiple formats, including plain text, tables, visualized charts, etc. Additionally, the assistant can complete tasks and return an action state. For transparency, responses can also be linked back to their original source documents, allowing users to trace how the information was derived.
@@ Acknowledgments
We thank the insightful discussions with Jie Fu, Jun Wang, Ying Wen, Zheng Tian, Qiuying Peng, and Grace Hui Yang.